The empirical formula of an unknown compound containing only carbon hydrogen and oxygen was found to be CH2O. In a separate experiment the molar mass of the molecule was determined to be 60 g/mol. What is the molecular formula of the compound?


Answer will be in C#H#O#


1
Expert's answer
2021-05-10T01:20:59-0400

Empirical formula = CH2O

Empirical formula mass =

12g/mol + 2g/mol + 16g/mol = 30g/mol


"\frac{Molecular formula mass}{Empirical formula mass}" = "\frac{60g/mol}{30g/mol}" = 2


Molecular formula = 2 × CH2O = C2H4O2